Circular No. 7590

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001W IN MCG +07-34-134 T. Puckett, M. Marcus, and D. George, Mountain Town, GA, report

the discovery of an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Feb. 26.46 UT. SN 2001W is located at α = 16^{h}41^{m}15^{s}.09, δ = +39^{o}17′44′′.5 (equinox 2000.0), which is 0′′.1 west and 20′′.6 north of the center of MCG +07-34-134. The new object was also present and confirmed on CCD frames taken on Feb. 28.34 with a 0.40-m reflector by J. Newton at Chiefland, FL. SN 2001W does not appear on unfiltered images taken by Puckett on 2000 June 29, July 8, and Aug. 23 or on Palomar Sky Survey images taken on 1993 Apr. 30, 1989 May 31 (limiting mag ∼ 21.0), and 1953 May 19 (limiting mag ∼ 19.7).

NOVA SAGITTARII 2001 Photographic magnitudes from T-Max 400 film, communicated in part

by T. Kato, Kyoto University: Feb. 2.843 UT, [10.0 (H. Nishimura, Kakegawa, Shizuoka, Japan); 17.832, [10.0 (Nishimura); 19.831, [13.7 (K. Takamizawa, Saki-machi, Nagano, Japan); 24.836, 8.1 (Nishimura). W. Liller, Viña del Mar, Chile, reports the following CCD broadband-V magnitudes: Feb. 25.367, 8.07; 26.335, 8.68; 27.374, 9.09; 28.341, 9.28. Visual magnitude estimates: Feb. 25.215, 8.5 (P. Schmeer, Bischmisheim, Germany); 25.63, 8.9 (A. Jones, Nelson, New Zealand); 26.250, 9.3 (A. Amorim, Florianopolis, Brazil); 26.835, 9.9 (A. Pearce, Nedlands, W. Australia); 27.205, 9.6 (A. Kammerer, Ettlingen, Germany); 27.794, 10.0 (Pearce); 28.635, 10.1 (J. Bedient, Honolulu, HI).

S/2001 (87) 1 A. Storrs, Towson University, and the Hubble Space Telescope (HST)

Asteroid Team (F. Vilas, R. Landis, E. Wells, C. Woods, B. Zellner, and M. Gaffey) confirm the companion to (87) Sylvia, reported on IAUC 7588, in HST images made on Feb. 23.57 UT. Six images show a companion at p.a. 306^{o} and separation 0′′.33. The companion appears bluer than does (87). Preliminary determinations of the magnitude difference between the two are as follows (listed are filter name, wavelength, and difference in magnitude): F439W, 439 nm, 6.00; F673N, 673, 6.27; F791W, 791, 6.44; F953N, 953, 6.67; F1042M, 1042, 9.37. The detections were marginal in the last two filters (the sixth image was saturated).

2001 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green